# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

### v2

@@ 4 Experiments
We conducted comparative experiments among PiSSA, LoRA, and full parameter fine-tuning on three models: LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50]. Our comparison spanned a diverse range of tasks, encompassing mathematics [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6]. PiSSA exhibited superior performance across these downstream tasks (Section 4.1). When combined with quantization techniques, PiSSA effectively reduces quantization errors. Compared to QLoRA [10] and LoftQ [47], PiSSA exhibits superior performance both during initialization and fine-tuning stages (Section 4.2). The ablation studies, which varied across different ranks, showed that PiSSA converged more swiftly and aligned more closely with the training data than LoRA, demonstrating robust superiority under equivalent trainable parameter configurations (Section 4.3). Further experiments demonstrated that employing the Fast SVD technique [56] allowed PiSSA to achieve a balance between initialization speed and performance (Section 4.4). Lastly, adapters initialized with principal singular values and vectors (PiSSA) were compared to those initialized with intermediate or less significant ones, revealing that direct fine-tuning of the model’s principal components yielded better outcomes (Section 4.5).
In our experiments, we adopt the Alpaca [57] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. We utilize the bf16 computation type, ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model. We train only on the first 100,000 conversions from the dataset for one epoch, and only using the response of the conversions to compute loss. The experiments in this section were conducted on the NVIDIA A800-SXM4(80G) GPU.
4.1 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
Table 1 presents a comparative evaluation of fine-tuning strategies using PiSSA, LoRA, and full parameter fine-tuning. In these experiments, we fine-tune LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [51] and MATH [2] validation sets. Additionally, the models are fine-tuned on the CodeFeedback dataset [53] and their coding proficiency is evaluated on the HumanEval [54] and MBPP [55] datasets. Furthermore, the models are trained on the WizardLM-Evol-Instruct dataset [7] and tested for their conversational abilities on the MT-Bench dataset [6].
Table 1: Performance of different models on various tasks
@@ 4.1 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
LLaMA 2-7B Full Fine-tuning 6738M 49.05 7.22 21.34 35.59 4.91
LoRA 320M 42.3 5.5 18.29 35.34 4.58
@@
Mistral-7B Full Fine-tuning 7242M 67.02 18.6 45.12 51.38 4.95
LoRA 168M 67.7 19.68 43.9 58.39 4.9
@@
Gemma-7B Full Fine-tuning 8538M 71.34 22.74 46.95 55.64 5.4
LoRA 200M 74.9 31.28 53.66 65.41 4.98
@@
As can be seen from Table 1, across all models and tasks, fine-tuning with PiSSA surpasses the effects of fine-tuning with LoRA, as well as most of full fine-tuning. In the future, we will extend the comparing to more models and tasks.
4.2 Comparative Analysis of QLoRA, LoftQ, and PiSSA in Quantization Fine-tuning
In this section, we compare the performance of QLoRA, LoftQ, and PiSSA in fine-tuning with quantization. All base models are quantized to ‘nf4’ precision and frozen, while the adapters are trained using ‘bfloat16’ precision. We begin by comparing the quantization errors of the three methods. For each method, the quantized baseline model is combined with its adapter, and the difference between the original weights and this combination serves as the error matrix. We measure the magnitude of the errors by the average nuclear norm of all error matrices. Using the quantization error of the original model without any adapter as a baseline, the errors for QLoRA, LoftQ, and PiSSA are divided by this baseline to express them as error rates. As QLoRA’s adapter employs Gaussian-zero initialization, the error rate does not decrease; LoftQ can eliminate quantization errors through multiple singular value decompositions, but for consistency with PiSSA, only one decomposition is performed. PiSSA requires a full-precision singular value decomposition, storing the principal singular values and vectors in the adapter before quantizing the residual model. We train LLaMA 2-7B on MetaMathQA and validate the effectiveness of the three methods on GSM8K and MATH datasets. The reductions in error and the training results are displayed in Figure 2.
Figure 2: Comparing the quantization error, the fine-tuning loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets.
The results, shown in Figure 2, reveal that PiSSA significantly outperforms LoftQ in reducing errors, achieving a 19% reduction at rank 128, as opposed to LoftQ’s 8%. LoftQ’s fine-tuning improvements over QLoRA are not significant, and in certain settings, it even underperforms compared to QLoRA. PiSSA effectively reduces quantization errors, and the training of principal singular values and vectors additionally provides benefits, thus substantially enhancing training outcomes in comparison to QLoRA and LoftQ.
4.3 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
@@
This section explores the impact of incrementally increasing the rank of PiSSA and LoRA from 1 to 128, aiming to determine the optimal number of trainable parameters needed to achieve performance comparable to full parameter training. The training is conducted using the MetaMathQA dataset, while the validation is performed on the GSM8K and MATH datasets. The outcomes of these experiments are depicted in Figure 3, with additional results presented in Figure 5 in the Appendix.
(a) Variation of loss with respect to rank 1 throughout the training phase. Additional ranks are depicted in Figure 5 in the Appendix.
(b) Final training loss across different ranks.
(c) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
(d) Rank-wise performance evaluated using pass@1 on the MATH dataset.
From Figure 3(a), it is observed that at the beginning of training, PiSSA quickly starts to converge, whereas LoRA experiences a period where the loss not only fails to decrease but even slightly increases. Furthermore, throughout the entire training process, the loss of PiSSA consistently remains at a lower level compared to LoRA. Figure 3(b) shows the final loss on the training set for models trained with ranks varying from 1 to 128. The data indicate that, compared to LoRA, PiSSA achieves a level of fitting to the training set similar to that of full-parameter fine-tuning with fewer trainable parameters. In Figures 3(c) and 3(d), we compare the performance of fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA outperforms LoRA with the same amount of trainable parameters. Furthermore, as the rank increases, PiSSA quickly reaches and surpasses the performance of full-parameter fine-tuning. Notably, when the training loss of LLaMA 2-7B with a rank of 128 PiSSA equals that of full-parameter fine-tuning, it surpasses full-parameter fine-tuning in terms of generalization on downstream tasks. This validates our approach of using principal singular values and vectors to replace full-parameter model fine-tuning, ignoring components of lower rank, resulting in a model with improved robustness.
4.4 Fast Singular Value Decomposition
@@
In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [56] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition. We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table 4.4. Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds. Initialization error indicates the magnitude of the discrepancy introduced by Fast SVD compared to SVD after decomposing the matrix. Specifically, the error is the sum of the absolute differences between the matrices decomposed by original SVD and Fast SVD. For the error, we report the results of the self-attention module in the table. Loss refers to the loss value at the end of training. In Fast SVD, the parameter niter refers to the number of subspace iterations to conduct. A larger niter leads to increased decomposition time but results in smaller decomposition error. The symbol $\infty$ represents the experimental results with the SVD method.
@@
4.5 Verifying the importance of principal singular values and vectors
5 Conclusion
This paper introduces a parameter-efficient fine-tuning (PEFT) technique in which singular value decomposition is applied to the weight matrix of pre-trained models to extract principal components. These components are then used to initialize an adapter named PiSSA. Fine-tuning PiSSA can closely replicate the effects of fine-tuning the complete model in the beginning, while maintaining a good parameter efficiency. Through extensive experiments, we show that PiSSA’s fine-tuning performance significantly surpasses that of LoRA. PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza. As PiSSA shares the same architecture as LoRA, it is easy to use and provides an efficient alternative initialization method. There remain several questions to be verified in the future: 1) Can PiSSA achieve significant improvements on a broader range of tasks and larger models? 2) When the iteration steps of LoRA are sufficiently long (adequately fitting the data), can it match the performance of PiSSA? 3) Can combining the successors of LoRA with PiSSA lead to further enhancement? 4) How to give a theoretical explanation for the advantages of PiSSA over LoRA? We are actively exploring these questions. Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments, and look forward to more tests and suggestions from the community.
Appendix A Equivalently Convert PiSSA into LoRA

### v3

@@ 5 Experiments
The experiments were conducted on the NVIDIA A800-SXM4(80G) GPU. In our experiments, we adopt the Alpaca [52] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. As discussed in Section B.3 of QLoRA [10], we compute the loss using only the responses from the instruction-following datasets. We ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model. We utilize the Float32 computation type for both the base model and the adapter in LoRA and PiSSA. For QLoRA, LoftQ, and QPiSSA, we use 4-bit NormalFloat [10] for the base model and Float32 for the adapter. BFloat16 [53] is used for full parameter fine-tuning to save the resources (see Appendix D).
5.1 Evaluating the Performance of PiSSA on both NLG and NLU Tasks
Table 1 presents a comparative evaluation of fine-tuning strategies PiSSA, LoRA, and full parameter fine-tuning on natural language generation (NLG) tasks. We fine-tuned LLaMA 2-7B [54], Mistral-7B-v0.1 [55], and Gemma-7B [56] on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [57] and MATH [2] validation sets. Additionally, the models were fine-tuned on the CodeFeedback dataset [58] and evaluated for coding proficiency using the HumanEval [59] and MBPP [60] datasets. Furthermore, the models were trained on the WizardLM-Evol-Instruct dataset [7] and tested for conversational abilities on the MT-Bench dataset [6]. All experiments were conducted using subsets containing 100K data points and were trained for only one epoch to reduce training overhead.
Table 1: Comparing PiSSA and LoRA on NLG tasks
@@ 5.1 Evaluating the Performance of PiSSA on both NLG and NLU Tasks
LLaMA 2-7B Full FT 6738M 49.05 7.22 21.34 35.59 4.91
LoRA 320M 42.30 5.50 18.29 35.34 4.58
@@
Mistral-7B Full FT 7242M 67.02 18.6 45.12 51.38 4.95
LoRA 168M 67.70 19.68 43.90 58.39 4.90
@@
Gemma-7B Full FT 8538M 71.34 22.74 46.95 55.64 5.40
LoRA 200M 74.90 31.28 53.66 65.41 4.98
@@
As shown in Table 1, across all models and tasks, fine-tuning with PiSSA consistently surpasses the performance of fine-tuning with LoRA. For instance, fine-tuning the LLaMA, Mistral, and Gemma with PiSSA results in performance improvements of 10.77%, 5.26%, and 3.04% respectively in mathematical tasks. In coding tasks, the improvements are 20%, 6.95%, and 1.14% respectively. In MT-Bench, we observe enhancements of 6.33%, 8.98%, and 13.25%. Notably, using PiSSA with only 2.3% of Gemma’s trainable parameters outperforms full parameter fine-tuning by 15.59% in coding tasks. Further experiments demonstrated that this improvement is robust across various amounts of training data and epochs (Section 5.2), including both 4-bit and full precision (Section 5.3), different model sizes and types (Section 5.4), and varying proportions of trainable parameters (Section 5.5).
Table 2: Comparing PiSSA and LoRA on NLU tasks. Details can be found in Appendix  I.
Method Parameters MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
RoBERTa-large (355M)
Full FT 355M 90.2 96.4 90.9 68.0 94.7 92.2 86.6 91.5
LoRA 1.84M 90.6 96.2 90.9 68.2 94.9 91.6 87.4 92.6
PiSSA 1.84M 90.7 96.7 91.9 69.0 95.1 91.6 91.0 92.9
DeBERTa-v3-base (184M)
Full FT 184M 89.90 95.63 89.46 69.19 94.03 92.40 83.75 91.60
LoRA 1.33M 90.65 94.95 89.95 69.82 93.87 91.99 85.20 91.60
PiSSA 1.33M 90.43 95.87 91.67 72.64 94.29 92.26 87.00 91.88
We also evaluate PiSSA’s natural language understanding (NLU) capability on the GLUE benchmark [61] with RoBERTa-large [62] and DeBERTa-v3-base [63]. Table 2 presents the results across 8 tasks conducted with the two base models. PiSSA outperforms LoRA on 14 out of 16 experimental settings, and achieves the same performance on QQP + RoBERTa. The single exception arises from the DeBERTa-based model on the MNLI dataset. Upon reviewing the training loss, we observed that PiSSA’s average loss of $0.17$ was lower than LoRA’s $0.24$ in the final epoch. This indicates that the fitting ability of PiSSA remains stronger than that of LoRA.
5.2 Experiments using Full Data and More Epochs
In this section, we finetune LLaMA 2-7B model on the complete MetaMathQA-395K dataset for 3 epochs to ensure thorough saturation. The training loss and gradient norms is visualized to demonstrate quicker convergence and evaluated on the GSM8K dataset every 1000 steps to demonstrate superior performance of PiSSA compared to LoRA. The results are depicted in Figure 4. Additionally, similar comparisons on Mistral-7B and Gemma-7B are detailed in Appendix G.
(a) Loss over steps.
(b) Grad norm over steps.
(c) Accuracy on GSM8K over training steps.
Figure 4: The loss, grad norm, and evaluation accuracy over the training steps of LoRA (indicated in blue), PiSSA (in orange), and full parameter fine-tuning (in green).
According to Figure 4(a), the loss of PiSSA reduces rapidly during the first 100 steps, and the grad norm (shown in Figure 4(b)) of PiSSA is significantly higher than that of LoRA, with a trend similar to full fine-tuning. Throughout the process, the loss of PiSSA remains lower than that of LoRA, indicating that PiSSA converges to a better local optimum. As shown in Figure 4(c), PiSSA consistently achieves higher accuracy compared to LoRA, and in most cases also surpasses full parameters fine-tuning. We hypothesize that this is because PiSSA is a denoised version of full fine-tuning. Comparing the grad norm and loss curves of PiSSA and full fine-tuning, we can see that the larger grad norm of full fine-tuning does not bring lower loss, indicating that a portion of the grad norm is spent on noisy directions not beneficial for loss reduction. This phenomenon is consistent with Figure 2(a).
5.3 Conducting 4-bit Quantization Experiments
In this section, we first compare the initial quantization error reduction ratio of PiSSA, QLoRA, and LoftQ. This ratio is defined as $(1-\frac{||W-(nf4(W^{{}^{\prime}})+AB)||_{*}}{||W-nf4(W)||_{*}})\times 100\%$, measuring the relative error decrease achieved by each mehod compared to directly quantizing the base model. The partial results are presented in Table 3, and the complete results can be found in Table 6 in Appendix E.
Table 3: The quantization error reduction ratio of QLoRA, LoftQ, and PiSSA across different layers.
Method Rank Q K V O Gate Up Down AVG
LLaMA 2-7B QLoRA 0 0 0 0 0 0 0 0
loftQ 128 16.5 16.5 15.9 16.0 12.4 12.4 12.3 14.6
PiSSA 128 27.9 27.2 18.7 18.6 15.8 13.6 13.6 19.4
LLaMA 3-8B QLoRA 0 0 0 0 0 0 0 0
loftQ 128 16.4 29.8 28.8 16.1 11.9 11.7 11.7 18.1
PiSSA 128 26.3 41.7 32.3 20.1 14.4 12.5 12.9 22.9
LLaMA 3-70B QLoRA 0 0 0 0 0 0 0 0
LoftQ 64 6.1 17.8 17.0 6.0 4.3 4.4 4.2 8.5
PiSSA 64 15.7 34.2 18.9 7.5 6.7 5.7 4.7 13.4
PiSSA 128 23.2 49.0 30.5 12.5 10.1 8.8 8.2 20.3
In Table 3, PiSSA reduces the quantization error by about 20% compared to directly quantizing the base model. The reduction is more significant for lower-rank matrices. For instance, in the LLaMA-3-70B [64], all “Key” projection layers see a reduction of 49%. The results in Table 3 validate that QLoRA, discussed in Section 4, does not reduce quantization error. In contrast, PiSSA significantly outperforms LoftQ in reducing quantization error, as further discussed in Appendix F.
Besides reducing the quantization error, we expect QPiSSA to also converge faster than QLoRA and LoftQ. We train LLaMA 3-8B with LoRA/QLoRA, PiSSA/QPiSSA, LoftQ, and full fine-tuning on MetaMathQA-395K for 3 epochs and record the loss, grad norm, and accuracy on GSM8K.
(a) Loss over training steps.
(b) Grad norm over training steps.
(c) Accuracy on GSM8K over training steps.
Figure 5: The loss, grad norm, and evaluation accuracy over the training steps of (Q)LoRA, (Q)PiSSA, LoftQ and full parameter fine-tuning.
According to Figure 5, QPiSSA’s loss reduction speed in the first 100 steps is even faster than PiSSA and full fine-tuning. Although LoftQ can reduce the quantization error, its loss convergence speed is not faster than LoRA and QLoRA, indicating that QPiSSA’s ability to reduce the quantization error and its fast convergence might also be orthogonal capabilities. After sufficient training, QPiSSA’s loss is also much lower than that of LoRA/QLoRA and LoftQ. The grad norm is significantly larger than those of LoRA/QLoRA and LoftQ. In terms of fine-tuning performance, QPiSSA’s accuracy is higher than that of QLoRA and LoftQ and even better than that of full-precision LoRA.
5.4 Experiments Across Various Sizes and Types of Models
In this section, we compare (Q)PiSSA and (Q)LoRA across 9 models, ranging from 7-70B parameters, including LLaMA 2-7/13B [54], LLaMA-3-8/70B [64], Mistral-7B [55], Gemma-7B [56], and Qwen1.5-7B [65], Yi-1.5-34B [66] and MoE models: DeepSeek-MoE-16B [67] and Mixtral-8x7B [68]. These models were fine-tuned on the MetaMathQA-100K and CodeFeedback-100K dataset and evaluated on the GSM8K and HumanEval. DeepSeek-MoE-16B, Mixtral-8x7B, Yi-1.5-34B, and LLaMA-3-70B were fine-tuned with QPiSSA and QLoRA, while the other models were using PiSSA and LoRA. From Figure 6, it can be observed that (Q)PiSSA, compared to (Q)LoRA, shows improved accuracy across various sizes and types of models, demonstrating its consistent advantage over (Q)LoRA.
Figure 6: Comparison of (Q)PiSSA and (Q)LoRA across models from 7B to 70B.
5.5 Experiments on Various Ranks
@@
This section explores the impact of incrementally increasing the rank of PiSSA/QPiSSA and LoRA/QLoRA from 1 to 128, aiming to determine whether PiSSA/QPiSSA consistently outperforms LoRA/QLoRA under different ranks. The training is conducted using the MetaMathQA-100K dataset for 1 epoch, while the validation is performed on the GSM8K and MATH datasets. The outcomes of these experiments are depicted in Figure 7, with additional results presented in Appendix H.
(a) Quantization error reduction ratio across ranks.
(b) Training loss under various ranks.
(c) Accuracy on GSM8K under various ranks.
(d) Accuracy on MATH under various ranks.
Figure 7: The comparison among (Q)LoRA, (Q)PiSSA, LoftQ, and full fine-tuning across ranks.
Figure 7(a) illustrates the quantization error reduction ratio across various ranks. In this figure, QLoRA shows no reduction in quantization error, while QPiSSA consistently outperforms LoftQ in reducing quantization error across all ranks, with a particularly notable advantage at lower ranks. In Figure 7(b), the final loss on the training set is shown for models trained with ranks ranging from 1 to 128. The results indicate that PiSSA and QPiSSA achieve a better fit to the training data compared to LoRA, QLoRA, and LoftQ. In Figures 7(c) and Figures 7(d), we compare the accuracy of the fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA consistently outperforms LoRA with the same amount of trainable parameters. Furthermore, as the rank increases, PiSSA will reach and surpass the performance of full-parameter fine-tuning.
6 Conclusion
This paper presents a PEFT technique that applies singular value decomposition (SVD) to the weight matrix of pre-trained models. The principal components obtained from the SVD are used to initialize a low-rank adapter named PiSSA, while the residual components are kept frozen, in order to achieve effective fine-tuning and parameter efficiency at the same time. Through extensive experiments, we found that PiSSA and its 4-bit quantization version QPiSSA significantly outperform LoRA and QLoRA in both NLG and NLU tasks, across different training steps, various model sizes and types, and under various amount of trainable parameters. PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza. As PiSSA shares the same architecture as LoRA, it can be seamlessly used in existing LoRA pipelines as an efficient alternative initialization method.
7 Limitation
There are still some questions with PiSSA not addressed in this paper: 1) Besides language models, can PiSSA also be adapted to convolutional layers and enhance the performance of vision tasks? 2) Can PiSSA also benefit from some improvements to LoRA, such as AdaLoRA [69] and DyLoRA [42] which adaptively adjust the rank? 3) Can we provide more theoretical explanations for the advantages of PiSSA over LoRA? We are actively exploring these questions. Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
Appendix A Conductive Experiments on Various SVD Components
Appendix B Fast Singular Value Decomposition
@@
In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [50] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition. We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table B. Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds. Initialization error indicates the magnitude of the discrepancy introduced by Fast SVD compared to SVD after decomposing the matrix. Specifically, the error is the sum of the absolute differences between the matrices decomposed by original SVD and Fast SVD. For the error, we report the results of the self-attention module in the table. Loss refers to the loss value at the end of training. In Fast SVD, the parameter niter refers to the number of subspace iterations to conduct. A larger niter leads to increased decomposition time but results in smaller decomposition error. The symbol $\infty$ represents the experimental results with the SVD method.
@@
Appendix C Equivalently Converting PiSSA into LoRA